# Weighted Aggregation of Conformity Scores for Classification

### v1

@@ Abstract
Conformal prediction is a powerful framework for constructing prediction sets with valid coverage guarantees in multi-class classification. However, existing methods often rely on a single score function, which can limit their efficiency and informativeness. We propose a novel approach that combines multiple score functions to improve the performance of conformal predictors by identifying optimal weights that minimize prediction set size. Our theoretical analysis establishes a connection between the weighted score functions and subgraph classes of functions studied in Vapnik-Chervonenkis theory, providing a rigorous mathematical basis for understanding the effectiveness of the proposed method. Experiments demonstrate that our approach consistently outperforms single-score conformal predictors while maintaining valid coverage, offering a principled and data-driven way to enhance the efficiency and practicality of conformal prediction in classification tasks.
@@ 1 Introduction
In the era of data-driven decision-making, the ability to quantify uncertainty in machine learning predictions is becoming increasingly important, particularly in high-stakes applications such as medical criminal justice (Berk et al., 2021) and diagnostics (Angelopoulos et al., 2024). While most machine learning methods focus on providing point predictions, supplementing these predictions with valid uncertainty estimates in the form of prediction sets or intervals, is crucial for responsible decision-making.
@@
Conformal prediction (Vovk et al., 2005; Manokhin, 2022) is a powerful framework that provides valid prediction sets with finite-sample coverage guarantees, regardless of the true underlying distribution (Angelopoulos et al., 2023). The key idea is to construct a prediction set for a new test point based on training data, such that the probability of the true label being contained in the prediction set is at least $1-\alpha$. This coverage guarantee holds irrespective of the point prediction algorithm used, making conformal prediction a versatile tool for quantifying uncertainty in machine learning predictions.
Central to conformal prediction is the concept of score functions, which quantify the conformity of a data point to the underlying distribution. In the split conformal prediction framework (Papadopoulos et al., 2002; Lei & Wasserman, 2014; Vovk et al., 2018), the training data is divided into a proper training set and a calibration set. The point prediction algorithm is trained on the proper training set, while the score function is evaluated on the calibration set. The conformal prediction set is then constructed by including all points whose score falls below a certain quantile of the calibration scores, with the quantile determined by the desired coverage level $1-\alpha$.
@@
The choice of score function plays a crucial role in determining the efficiency of the resulting prediction sets, with well-chosen score functions leading to more informative and precise predictions. The flexibility in choosing score functions allows conformal prediction to adapt to the specific characteristics of the data and the underlying distribution. This is an active area of research in conformal prediction, with ongoing efforts to develop score functions that yield the most informative and efficient prediction sets for a given problem setting, including regression (Papadopoulos et al., 2008, 2011; Romano et al., 2019; Kivaranovic et al., 2020; Guan, 2023; Colombo, 2023, 2024) and classification (Sadinle et al., 2019; Romano et al., 2020; Angelopoulos et al., 2021; Huang et al., 2024; Luo & Zhou, 2024). Despite the dependence of prediction set efficiency on the choice of score function and the performance of the underlying point prediction algorithm, the coverage guarantee of conformal prediction remains intact, making it a robust tool for providing reliable uncertainty estimates in machine learning predictions.
In this work, we propose a novel approach to optimize a linear combination of conformity scores using efficiency as the objective function. Our goal is to search for optimal weights for such linear combinations, enabling us to fully utilize the advantages of existing score functions. While our approach shares similarities with Izbicki et al. (2019) and Yang & Kuchibhotla (2024), which discuss the construction of the smallest prediction set using kernel density estimators and families of machine learning algorithms, respectively, our method distinguishes itself in three key aspects:
@@
    First, our approach emphasizes finding an optimal combination of score functions, which is more practical in the current landscape of numerous high-quality and diverse score functions, rather than focusing on training multiple algorithms. By leveraging the strengths of existing score functions, we aim to improve the efficiency of conformal prediction while maintaining its coverage guarantees.
@@
    Second, we investigate and categorize major data split strategies for determining optimal weights for combining score functions. These strategies include Validity First Conformal Prediction (VFCP) and Efficiency First Conformal Prediction (EFCP) (Yang & Kuchibhotla, 2024), Data Leakage Conformal Prediction (DLCP), and its variant that utilizes all available data (DLCP+). We establish the coverage property for VFCP and quantify the coverage gap for the other three strategies, along with their corresponding efficiency improvements.
@@
    Third, our method distinguishes itself from selection methods that focus on identifying the single best-performing score function. By aggregating multiple score functions through an optimally chosen weight, our approach can achieve performance levels that surpass those of any individual score function. This ability to harness the collective strengths of multiple score functions sets our method apart and enables it to provide more efficient and informative prediction sets.
Our weighted aggregation approach addresses the challenge of constructing efficient and valid prediction sets in multi-class classification by learning optimal weights for combining different score functions using various data split procedure. This allows for adaptation to the specific characteristics of the dataset and potentially outperforms methods relying on a single score function. We provide a theoretical analysis of our approach, establishing finite-sample guarantees for coverage and width and deriving oracle inequalities to quantify its performance relative to an “optimal” aggregation. Through extensive experiments, we demonstrate the superior performance of our weighted aggregation approach compared to single-score methods, highlighting its potential for improving the efficiency of conformal prediction in classification tasks.
@@
The rest of the paper is organized as follows. In Section 2, we present our weighted aggregation approach and the various data split strategies. Section 3 contains our theoretical analysis, establishing coverage guarantees and deriving the expected prediction set sizes. In Section 4, we discuss our experimental results, demonstrating the effectiveness of our approach. Related works are presented in Section 5. Finally, we conclude the paper in Section 6 and discuss future research directions.
2 Weighted Aggregation of Score Functions
2.1 Basic Conformal Prediction for Classification
@@
We start by assuming that a classification algorithm provides $\widehat{p}_{y}(x)$, which approximates $P(Y=y|X=x)$. While our method and theoretical analysis do not depend on the accuracy of this approximation, it is beneficial to assume that higher values of $\widehat{p}_{y}(x)$ indicate a greater likelihood of sample $x$ having label $y$. We consider this training procedure to be performed on a separate dataset, ensuring that $\widehat{p}(y|x)$ is independent of the dataset used in this paper.
@@
Let us first consider a single non-conformity score function $s(x,y)$. This function is defined such that higher values of $s(x,y)$ indicate a higher priority of believing $x$ has label $y$. A common choice for the non-conformity score is $s(x,y)=\widehat{p}_{y}(x)$ Sadinle et al. (2019). In this context, conformal prediction for classification problems can be described by the following algorithm. To summarize, we first find a threshold in a set of labeled data, so that $s(x_{i},y_{i})\geq Q_{1-\alpha}$ holds for at least $1-\alpha$ proportion in the set $\mathcal{I}_{\textrm{train}}$. Then we use this threshold to define the prediction set for $x_{i}$ in the test set, which is the upper level set of the function $y\mapsto s(x_{i},y)$.
@@
This algorithm constructs prediction sets $\widehat{C}_{1-\alpha}(x_{i})$ for each $i\in\mathcal{I}_{\textrm{test}}$, based on the non-conformity scores and a threshold determined by the desired coverage probability $1-\alpha$. The algorithm splits the training data into two subsets: $\mathcal{I}_{1}$ for calculating the non-conformity scores, and $\mathcal{I}_{2}$ for calibration. The threshold $Q_{1-\alpha}$ is chosen to ensure the desired coverage probability.
A key property of this method is that, under the assumption of exchangeability of the data points (a weaker assumption than i.i.d.), it guarantees that the resulting prediction sets will contain the true label with probability at least $1-\alpha$. This property holds regardless of the accuracy of the underlying classification algorithm, making conformal prediction a powerful tool for uncertainty quantification.
2.2 Example of Score Functions for Classifications
In the context of conformal prediction, the choice of score function $s(x,y)$ is crucial as it directly impacts the resulting prediction sets. These score functions serve as non-conformity measures, quantifying how well a label fits a given input. Let $K$ denote the number of classes in our classification problem.
Various score functions have been proposed and their effectiveness demonstrated through extensive experimentation on diverse datasets. Three typical choices of score functions are:
@@ 2.2 Example of Score Functions for Classifications
    Threshold (THR) (Sadinle et al., 2019): $s_{1}(x,y)=\widehat{p}_{y}(x)$.
@@
    Adaptive Prediction Sets (APS) (Romano et al., 2020): $s_{2}(x,y)=\sum_{y^{\prime}\in[K]}\widehat{p}_{y^{\prime}}(x)\mathbf{1}\{\widehat{p}_{y^{\prime}}(x)\leq\widehat{p}_{y}(x)\}$.
@@
    Rank-based (RANK) (Luo & Zhou, 2024): $s_{3}(x,y)=|\{y^{\prime}\in[K]:\widehat{p}_{y^{\prime}}(x)<\widehat{p}_{y}(x)\}|$.
We can easily plug the APS and RANK score functions into Algorithm 1 to create prediction sets. While the original publications introducing these score functions presented additional complexities to guarantee precise $1-\alpha$ expected coverage, these adjustments do not significantly impact the presentation of our proposed approach and are therefore omitted for clarity. The reasoning behind these scores will not affect how we combine them or our theoretical analysis. In the following sections, we will demonstrate how these score functions can be effectively utilized in our proposed approach without the need for additional complexity to ensure exact coverage. A detailed introduction and discussion of these score functions, along with their adjustments, will be provided in Section 4.
2.3 Aggregation of Score Functions
In Yang & Kuchibhotla (2024), the authors propose aggregating score functions for conformal prediction. Their approach selects the score function that yields the smallest average prediction set size among the available score functions. They also introduce two data splitting methods: Efficiency First Conformal Prediction (EFCP) and Validity First Conformal Prediction (VFCP).
Given a vector of score functions $s(x,y)=(s_{1}(x,y),\dots,s_{d}(x,y))^{\top}$, instead of simply choosing the score that provides the smallest prediction set on a calibration set, we propose assigning weights to the scores and defining a new weighted score function:
@@ 2.3 Aggregation of Score Functions
2.4 Algorithm for Determining the Weight and the Threshold
  1. 1.
    Extract $\mathcal{I}_{1},\mathcal{I}_{2}$, and $\mathcal{I}_{3}$ from the dataset. In the meta version of the algorithm, these sets can be arbitrary, provided they are sufficiently large. However, for clarity and ease of understanding for first-time readers, it is helpful to assume that $\mathcal{I}_{1}=\mathcal{I}_{2}$ and that $\mathcal{I}_{1}$ and $\mathcal{I}_{3}$ are partitions of $\mathcal{I}_{\textrm{train}}$.
  2. 2.
    Define a finite subset $\mathcal{W}\subseteq\mathbb{R}^{d}$ for candidates of $\bm{w}$. Theoretically, $\mathcal{W}$ can be an arbitrary subset of $\mathbb{R}^{d}$. Given that the scores are all non-conformity scores, it is sometimes desirable for the weights to be non-negative. In this case, letting $\mathcal{W}$ be the grid of the simplex $\Delta(d)\subseteq\mathbb{R}^{d}$ is a typical choice. The compactness of $\Delta(d)$ can make the algorithm computationally feasible.
  3. 3.
    $Q_{1-\alpha}^{(1)}(\bm{w})=\lceil(1+|\mathcal{I}_{1}|)(1-\alpha)\rceil\text{-th smallest value of }\{\langle\bm{w},s(x_{i},y_{i})\rangle:i\in\mathcal{I}_{1}\}.$ (2)
    This corresponds to Step 5 in Algorithm 1.
  4. 4.
    $\widehat{C}_{1-\alpha}(x_{i},\bm{w},Q_{1-\alpha}^{(1)}(\bm{w}))=\{y\in[K]:\langle\bm{w},s(x_{i},y)\rangle\geq Q_{1-\alpha}^{(1)}(\bm{w})\}\quad\text{for}\quad i\in\mathcal{I}_{2}.$ (3)
  5. 5.
    Find the optimal weight vector $\widehat{\bm{w}}$ that minimizes the average prediction set size:
    $\widehat{\bm{w}}\in\arg\min_{\bm{w}\in\mathcal{W}}\sum_{i\in\mathcal{I}_{2}}|\widehat{C}_{1-\alpha}(x_{i},\bm{w},Q_{1-\alpha}^{(1)}(\bm{w}))|.$ (4)
  6. 6.
    $Q_{1-\alpha}^{(2)}=\lceil(1+|\mathcal{I}_{3}|)(1-\alpha)\rceil\text{-th smallest value of }\{\langle\widehat{\bm{w}},s(x_{i},y_{i})\rangle:i\in\mathcal{I}_{3}\}.$ (5)
  7. 7.
    The final output is the confidence set:
    $\widehat{C}_{1-\alpha}(x_{i},\widehat{\bm{w}},Q_{1-\alpha}^{(2)})=\{y\in[K]:\langle\widehat{\bm{w}},s(x_{i},y)\rangle\geq Q_{1-\alpha}^{(2)}\}\quad\text{for}\quad i\in\mathcal{I}_{\textrm{test}}.$ (6)
The following two key observations of this procedure:
  • (a)
    Once the optimal weight vector $\widehat{\bm{w}}$ has been determined, the subsequent steps of the calibration procedure, as described in Eq. (5), and the construction of the prediction set, as outlined in Eq. (6), are identical to Step 5 and Step 6 of Algorithm 1. The additional steps preceding these are designed to identify an appropriate weight vector $\bm{w}$ that optimizes the performance of the algorithm.
  • (b)
    The operation in Eq. (2) and Eq. (5) are the steps for finding quantiles, and Eq. (3) and Eq. (6) are the steps for defining prediction sets. For finding the quantiles, we need the labels of the samples, while for defining the prediction set, only the feature $x$ of the sample is required. In other words, it is possible to include the test set in Eq. (3) and Eq. (6).
In the following algorithm, we present a formalized version of the procedure outlined above.
@@ 2.4 Algorithm for Determining the Weight and the Threshold
1:$\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{\textrm{train}}}$: Labeled training data; $\{x_{i}\}_{i\in\mathcal{I}_{\textrm{test}}}$: Unlabeled test data; $\alpha\in(0,1)$: Pre-defined significance level; $\mathcal{W}\subseteq\mathbb{R}^{d}$: Set of possible weight vectors, where $d$ is the number of score functions.
2:$\big{(}C_{1-\alpha}(X_{i})\big{)}_{i\in\mathcal{I}_{\textrm{test}}}$: Sequence of prediction sets for test data with anticipated $1-\alpha$ coverage
@@
10:     $\big{(}\widehat{C}_{1-\alpha}(x_{i};\bm{w},Q_{1-\alpha}^{(1)}(\bm{w}))\big{)}_{i\in\mathcal{I}_{2}}\leftarrow$ Evaluation($\bm{w},\mathcal{I}_{2},Q_{1-\alpha}^{(1)}(\bm{w})$).
11:     Compute average prediction set size $S{(\bm{w})}=\frac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}|\widehat{C}_{1-\alpha}(x_{i};\bm{w},Q_{1-\alpha}^{(1)}(\bm{w})\big{)}|$.
12:end for
13:$\widehat{\bm{w}}\leftarrow\arg\min_{\bm{w}\in\mathcal{W}}S{(\bm{w})}$.
14:Step 3: Test-time Prediction Set Construction
15:$Q_{1-\alpha}^{(2)}\leftarrow$ Calibration($\widehat{\bm{w}},\mathcal{I}_{3},\alpha$).
16:$\big{(}\widehat{C}_{1-\alpha}(x_{i},\widehat{\bm{w}},Q_{1-\alpha}^{(2)})\big{)}_{i\in\mathcal{I}_{\textrm{test}}}\leftarrow$ Evaluation($\widehat{\bm{w}},\mathcal{I}_{\textrm{test}},\mathcal{I}_{3}$).
17:return $\big{(}\widehat{C}_{1-\alpha}(x_{i},\widehat{\bm{w}},Q_{1-\alpha}^{(2)})\big{)}_{i\in\mathcal{I}_{\textrm{test}}}$

### v2

# Averaging of Conformity Scores for Classification

@@ Abstract
Conformal prediction provides a robust framework for generating prediction sets with finite-sample coverage guarantees, independent of the underlying data distribution. However, existing methods typically rely on a single conformity score function, which can limit the efficiency and informativeness of the prediction sets. In this paper, we present a novel approach that enhances conformal prediction for multi-class classification by optimally averaging multiple conformity score functions. Our method involves assigning weights to different score functions and employing various data splitting strategies. Additionally, our approach bridges concepts from conformal prediction and model averaging, offering a more flexible and efficient tool for uncertainty quantification in classification tasks. We provide a comprehensive theoretical analysis grounded in Vapnik–Chervonenkis (VC) theory, establishing finite-sample coverage guarantees and demonstrating the efficiency of our method. Empirical evaluations on benchmark datasets show that our weighted averaging approach consistently outperforms single-score methods by producing smaller prediction sets without sacrificing coverage.
@@ 1 Introduction
Conformal prediction Vovk et al. (2005); Manokhin (2022) is a robust framework that generates prediction sets with finite-sample coverage guarantees, irrespective of the underlying data distribution Angelopoulos et al. (2023). The fundamental principle of conformal prediction is to construct a prediction set for a new test instance based on the training data, ensuring that the true label is included with a probability of at least $1-\alpha$. This coverage assurance holds regardless of the specific point prediction algorithm employed, making conformal prediction a versatile tool for uncertainty quantification in machine learning. In the split conformal prediction framework Papadopoulos et al. (2002); Lei and Wasserman (2014); Vovk et al. (2018), the training data is partitioned into a training set and a calibration set. The predictive model is trained on the training subset, while the score functions are evaluated on the calibration subset. The conformal prediction set then comprises all labels whose scores fall below a specific quantile of the calibration scores, with the quantile determined by the coverage level.
@@
The choice of score function is critical in determining the efficiency of the resulting prediction sets, especially for multi-class classification. Well-chosen score functions can lead to more informative and precise predictions. This flexibility allows conformal prediction to adapt to the specific characteristics of the data and distribution. Consequently, developing score functions that optimize informativeness and efficiency for various problem settings, including regression Papadopoulos et al. (2008, 2011); Romano et al. (2019); Kivaranovic et al. (2020); Guan (2023); Colombo (2023, 2024) and multi-class classification Sadinle et al. (2019); Romano et al. (2020); Angelopoulos et al. (2021); Huang et al. (2024); Luo and Zhou (2024), remains an active area of research. This work focuses on enhancing the efficiency of prediction sets for classification tasks. While the underlying idea can extend to regression, this paper will focus on the tasks of multi-class classification.
Our approach assumes the availability of multiple score functions for the same classification task, each differing due to variations in the classification algorithm or the definition of the score. We propose assigning optimal weights to aggregate these score functions. Using a validation set, we determine a threshold to achieve the desired coverage and identify the weight combination that minimizes the prediction set size. The final prediction is then based on this weighted score function. Our aim is to find the optimal weights for linear combinations of score functions, thereby fully leveraging the strengths of existing score functions. While our approach shares similarities with Yang and Kuchibhotla (2024), it stands out in three key aspects:
@@
    Weighted Averaging of Score Functions: Instead of selecting the single best-performing score function, our approach combines multiple score functions through optimal weighting. This averaging can yield more efficient prediction sets than any individual score function while maintaining the desired coverage guarantees.
@@
    Novel Data Splitting Strategies: We explore and categorize several data splitting methods to determine the optimal weights for combining score functions. In addition to Validity First Conformal Prediction (VFCP) and Efficiency First Conformal Prediction (EFCP) discussed in Yang and Kuchibhotla (2024), we introduce Data Leakage Conformal Prediction (DLCP) and its variant DLCP+, which utilize all available data to enhance weight determination.
@@
    Theoretical Foundations Using VC Theory: We provide a theoretical analysis of our method that leverages Vapnik–Chervonenkis theory to establish coverage guarantees and expected prediction set sizes. This solid mathematical foundation underscores the validity and efficiency of our approach across different data splitting strategies.
Beyond its novel contributions to the conformal prediction literature, our method is closely related to model averaging Claeskens and Hjort (2008), a well-established technique in machine learning. Unlike traditional model averaging, which assigns weights to different models to improve prediction accuracy, our method assigns weights to score functions. This distinction requires the development of specific data splitting techniques to ensure the desired coverage guarantees. Consequently, our work can be viewed as an innovative adaptation of model averaging principles to the conformal prediction framework.
@@
The remainder of the paper is organized as follows. In Section 3, we detail our weighted averaging approach and the various data splitting strategies employed. Section 3 presents the theoretical analysis, establishing coverage guarantees and expected prediction set sizes. In Section 4, we demonstrate the effectiveness of our method through experiments. Related works are discussed in Section 5. We conclude in Section 6 and outline future research directions.
2 Methodology
2.1 Conformal Prediction for Classification
@@
We start by assuming that a $K$-class classification algorithm provides $\widehat{p}_{y}(x)$, which approximates $P(Y=y|X=x)$ for $y\in[K]$. While our method and theoretical analysis do not depend on the accuracy of this approximation, it is beneficial to assume that higher values of $\widehat{p}_{y}(x)$ indicate a greater likelihood of sample $x$ having label $y$. We consider this training procedure to be performed on a separate dataset, ensuring that $\widehat{p}_{y}(x)$ is independent of the dataset used in this paper.
@@
Let us first consider a single non-conformity score function $s(x,y)$. This function is defined such that higher values of $s(x,y)$ indicate a higher priority of believing $x$ has label $y$. A common choice for the non-conformity score is $s(x,y)=\widehat{p}_{y}(x)$ Sadinle et al. (2019). In this context, conformal prediction for classification problems can be described by the following algorithm. To summarize, we first find a threshold in a set of labeled data, so that $s(x_{i},y_{i})\geq Q_{1-\alpha}$ holds for at least $1-\alpha$ proportion in the set $\mathcal{I}_{2}$. Then we use this threshold to define the prediction set for $x_{i}$ in the test set, which is the upper level set of the function $y\mapsto s(x_{i},y)$.
This algorithm constructs prediction sets $\widehat{C}_{1-\alpha}(x_{i})$ for each $i\in\mathcal{I}_{\textrm{test}}$, based on the non-conformity scores and a threshold determined by the desired coverage probability $1-\alpha$. The algorithm splits the training data into two subsets: $\mathcal{I}_{1}$ for calculating the non-conformity scores, and $\mathcal{I}_{2}$ for calibration. The threshold $Q_{1-\alpha}$ is chosen to ensure the desired coverage probability under the exchangeability assumption.
@@
$\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{\text{train}}}$$\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{3}}$$\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{1}}$$\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{2}}$ $Q_{1-\alpha}^{(1)}(\bm{w})$ by Eq. (2) $\big{(}\widehat{C}_{1-\alpha}(x_{i},\bm{w},Q_{1-\alpha}^{(1)}(\bm{w}))\big{)}_{i\in\mathcal{I}_{2}}$ by Eq. (3) $\widehat{\bm{w}}$ by Eq. (4) $Q_{1-\alpha}^{(2)}$ by Eq. (5) $\{x_{i}\}_{i\in\mathcal{I}_{\text{test}}}$ $\left(\widehat{C}_{1-\alpha}(x_{i},\widehat{\bm{w}},Q_{1-\alpha}^{(2)})\right)_{i\in\mathcal{I}_{\text{test}}}$ by Eq. (6)
Figure 1: This example illustrates a framework for data splitting into $\mathcal{I}_{1},\mathcal{I}_{2},\mathcal{I}_{3}$, and $\mathcal{I}_{\text{test}}$. Algorithm 2 presents the complete procedure. Briefly, $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ are used in Steps 1-2 to select the optimal weight $\widehat{\mathbf{w}}$, while $\mathcal{I}_{3}$ is used in Step 3 as the calibration set for $\mathcal{I}_{\text{test}}$ predictions. We present four options: VFCP, EFCP, DLCP, and DLCP+. Their coverage and size properties are discussed theoretically in Section 3 and empirically in Section 4.
2.2 Various Score Functions for Classifications
In conformal prediction, the choice of score function $s(x,y)$ critically influences the prediction sets. These functions act as non-conformity measures that quantify how appropriately a label matches an input.
Score functions are determined by two primary factors: (1) The predictive model’s quality in estimating $\widehat{p}_{y}(x)\approx\mathbb{P}(Y=y|X=x)$, and (2) The specific methodology for converting these probability estimates into scores. We will consider approaches:
@@ 2.2 Various Score Functions for Classifications
    Threshold (THR) Sadinle et al. (2019): $s_{\text{THR}}(x,y)=\widehat{p}_{y}(x)$
@@
    Adaptive Prediction Sets (APS) Romano et al. (2020): $s_{\text{APS}}(x,y)=\sum_{y^{\prime}\in[K]}\widehat{p}_{y^{\prime}}(x)\mathbf{1}\{\widehat{p}_{y^{\prime}}(x)\leq\widehat{p}_{y}(x)\}$
@@
    Rank-based (RANK) Luo and Zhou (2024): $s_{\text{RANK}}(x,y)=|\{y^{\prime}\in[K]:\widehat{p}_{y^{\prime}}(x)<\widehat{p}_{y}(x)\}|$
These score functions integrate directly with Algorithm 1. While the original formulations include additional refinements to achieve exact $1-\alpha$ coverage guarantees, these refinements do not materially affect our combination methodology and are omitted for simplicity. Our theoretical analysis remains invariant to specific score function design choices. Other notable approaches like RAPS Angelopoulos et al. (2021) and SAPS Huang et al. (2024) demonstrate alternative methodologies for deriving scores from $\widehat{p}_{y}(x)$.
2.3 Averaging Score Functions
In Yang and Kuchibhotla (2024), the authors propose their method with multiple score functions. Their approach selects the score function that yields the smallest average prediction set size among the available score functions. They also introduce two data splitting methods: Efficiency First Conformal Prediction (EFCP) and Validity First Conformal Prediction (VFCP).
Given a vector of score functions $s(x,y)=(s_{1}(x,y),\dots,s_{d}(x,y))^{\top}$, instead of simply choosing the score that provides the smallest prediction set on a calibration set, we propose assigning weights $\bm{w}\in\mathcal{W}\subseteq\mathbb{R}^{d}$ to the non-confomity scores and defining the weighted score function:
@@ 2.3 Averaging Score Functions
1:$\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{\textrm{train}}}$: Labeled training data; $\{x_{i}\}_{i\in\mathcal{I}_{\textrm{test}}}$: Unlabeled test data; $\alpha\in(0,1)$: Pre-defined significance level; $s_{i}(\cdot,\cdot),i=1,\dots,d$: score functions. $\mathcal{W}\subseteq\mathbb{R}^{d}$: Set of possible weight vectors.
2:$\big{(}C_{1-\alpha}(X_{i})\big{)}_{i\in\mathcal{I}_{\textrm{test}}}$: Sequence of prediction sets for test data with anticipated $1-\alpha$ coverage.
@@
10:     $\big{(}\widehat{C}_{1-\alpha}(x_{i};\bm{w},Q_{1-\alpha}^{(1)}(\bm{w}))\big{)}_{i\in\mathcal{I}_{2}}$
11:             $\leftarrow$ Evaluation($\bm{w},\mathcal{I}_{2},Q_{1-\alpha}^{(1)}(\bm{w})$).
12:     Compute average prediction set size $S{(\bm{w})}=\frac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}|\widehat{C}_{1-\alpha}(x_{i};\bm{w},Q_{1-\alpha}^{(1)}(\bm{w})\big{)}|$.
13:end for
14:$\widehat{\bm{w}}\leftarrow\arg\min_{\bm{w}\in\mathcal{W}}S{(\bm{w})}$.
15:Step 3: Test-time Prediction Set
16:$Q_{1-\alpha}^{(2)}\leftarrow$ Calibration($\widehat{\bm{w}},\mathcal{I}_{3},\alpha$).
17:$\big{(}\widehat{C}_{1-\alpha}(x_{i},\widehat{\bm{w}},Q_{1-\alpha}^{(2)})\big{)}_{i\in\mathcal{I}_{\textrm{test}}}$leftarrow Evaluation($\widehat{\bm{w}},\mathcal{I}_{\textrm{test}},\mathcal{I}_{3}$).
18:return $\big{(}\widehat{C}_{1-\alpha}(x_{i},\widehat{\bm{w}},Q_{1-\alpha}^{(2)})\big{)}_{i\in\mathcal{I}_{\textrm{test}}}$
2.4 The Optimal Weight and the Threshold
  1. 1.
    Extract $\mathcal{I}_{1},\mathcal{I}_{2},\mathcal{I}_{3}$ from dataset. Assume $\mathcal{I}_{1}=\mathcal{I}_{2}$ and $\mathcal{I}_{1},\mathcal{I}_{3}$ partition $\mathcal{I}_{\textrm{train}}$ for clarity.
  2. 2.
    Define $\mathcal{W}\subseteq\mathbb{R}^{d}$ for $\bm{w}$ candidates. For non-negative weights, use grid of simplex $\Delta(d)\subseteq\mathbb{R}^{d}$.
  3. 3.
    $Q_{1-\alpha}^{(1)}(\bm{w})=\ \lceil(1+|\mathcal{I}_{1}|)(1-\alpha)\rceil\text{-th smallest}\text{value of}\quad\{\langle\bm{w},s(x_{i},y_{i})\rangle:i\in\mathcal{I}_{1}\}.$ (2)
    This corresponds to Step 5 in Algorithm 1.
  4. 4.
    $\widehat{C}_{1-\alpha}(x_{i},\bm{w},Q_{1-\alpha}^{(1)}(\bm{w}))=\{y\in[K]:\langle\bm{w},s(x_{i},y)\rangle\geq Q_{1-\alpha}^{(1)}(\bm{w})\}.$ (3)
  5. 5.
    The optimal weight vector $\widehat{\bm{w}}$ is obtained by minimizing the empirical prediction set size:
    $\widehat{\bm{w}}\in\operatorname*{arg\,min}_{\bm{w}\in\mathcal{W}}\sum_{i\in\mathcal{I}_{2}}|\widehat{C}_{1-\alpha}(x_{i},\bm{w},Q_{1-\alpha}^{(1)}(\bm{w}))|,$ (4)
    where $\mathcal{W}$ is a discrete grid over the $(d-1)$-dimensional probability simplex $\Delta^{d-1}$. We systematically explore candidate weights via grid search with step size $\varepsilon=0.01$, as detailed in Appendix A.
  6. 6.
[…]
    value of $\{\langle\widehat{\bm{w}},s(x_{i},y_{i})\rangle:i\in\mathcal{I}_{3}\}.$
  7. 7.
    The final output is the confidence set for samples $i\in\mathcal{I}_{\textrm{test}}$ in the test set:
    $\widehat{C}_{1-\alpha}(x_{i},\widehat{\bm{w}},Q_{1-\alpha}^{(2)})=\{y\in[K]:\langle\widehat{\bm{w}},s(x_{i},y)\rangle\geq Q_{1-\alpha}^{(2)}\}.$ (6)
This procedure comprises two primary steps: Threshold Calibration: Determining appropriate thresholds using Equations Eq. (2) and Eq. (5) based on the calibration subsets $\mathcal{I}_{1}$ and $\mathcal{I}_{3}$. Prediction Set Evaluation: Constructing the prediction sets for $\mathcal{I}_{2}$ and $\mathcal{I}_{\text{test}}$ as outlined in Equations Eq. (3) and Eq. (6). These steps are encapsulated in Function 1 and Function 2, respectively. The entire procedure of the proposed method is summarized in Algorithm 2.